5. When the price of Ford pickup trucks rises from $18,000 to $19,000, the quantity of Chevy trucks demanded increases from 112,000 to 144,000. Using the midpoint method, what is the cross elasticity of demand between Ford and Chevy trucks?
Answer #1

Cross price elasticity = % change in the quantity demand for good X / % change in the price of good Y

%change in price = 18,000 - 19,000/ (18,000 + 19,000) / 2

= -1.35

% change in demand = (114,000 - 112000) / ((114,000 + 112000))/ 2

= -6.25

Cross price elasticity = 6.25 / 1.35 = 4.65

As the value is in positive these goods are substitute and the value of cross price elasticity is 4.65.